The Death Star (m =2.5x10^18kg) orbits the planet Yanet at a distance d = 220 km above the planet’s surface. Yanet has a radius of 3.4x10^4 m. Given that

the speed of the Death Star is v =6x10^3

m/s, what is the mass of the planet?
Answer #1

Let the mass of planet be M

Radius of planet=3.4*10^4 m=34 km

Distance of star from center of planet=220+34=254 km=254000 m

Velocity of star=6*10^3 m/s

Gravitational constant=6.67*10^-11

From equating centripetal force and gravitational Force

6.67*10^-11*M/254000=(6*10³)²

M=1.37*10^23 kg

This is the required answer